OpenStudy (anonymous):

Solve the linear equation. 2(x + 6) + 6 = 3(x + 5) + 7 A. {6} B. {-4} C. {12} D. {9}

OpenStudy (kira_yamato):

2(x + 6) + 6 = 3(x + 5) + 7 2x + 12 + 6 = 3x + 15 + 7 18 - 22 = x x = -4 ==> (B)